OpenStudy (anonymous):

a card is drawn from a standard deck of cards. determine whether the events are mutually exclusive or inclusive. then, find the probability: P(a 4 or a club)

OpenStudy (anonymous):

is there a 4 of clubs? if so then not mutually exclusive

OpenStudy (anonymous):

13 clubs and 4 "fours" but you don't count the 4 of clubs twice, so answer is 16/52

OpenStudy (amistre64):

P(4 or club) = P(4) + P(club) - P(4andClub)

OpenStudy (anonymous):

\[P(A\cup B)=P(A)+P(B)-P(A \cap B)\]

OpenStudy (amistre64):

\[\frac{4}{52}+\frac{13}{52}-\frac{1}{52}\]

OpenStudy (anonymous):

true, but i like to count. especially cards and dice

OpenStudy (amistre64):

257 = 1025 .... ☺

OpenStudy (amistre64):

☺=☺

OpenStudy (anonymous):

is it inclusive or exclusive